# Finite-Time Behavior of Erlang-C Model: Mixing Time, Mean Queue Length and Tail Bounds

### v3

@@ 5.1. Proof of Lemma 2
(37) $\tau(q)$ $\geq 3n^{1-\alpha}\frac{13}{80}\frac{e^{\frac{-1}{12}}\sqrt{3}n^{\alpha-1}}{4}=\frac{39\sqrt{3}e^{\frac{-1}{12}}}{320}\text{ for }q=n-1,$
(38) $\tau(q)$ $\geq 2e^{2}n^{1-2\alpha}\times\frac{13}{80}\frac{e^{\frac{-1}{12}}\sqrt{3}n^{\alpha-1}}{4}=\frac{13\sqrt{3}e^{\frac{23}{12}}}{160}n^{-\alpha}\text{ for }q\in K-\{n-1\}.$
@@
$T_{1}\leq 560229\times n^{2-4\alpha}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$.
@@
$T_{2}\leq 459951\times n^{3-6\alpha}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1).$
Finally, we show that the total mass of $\tau$ is no more than a constant factor of the total mass of $\nu_{K}$.
Claim 3.
$\sum_{q\in K}\tau(q)\leq 340$ for $K=\{\lceil 2\lambda_{n}\rceil-n,...,n-1\}$.
@@
The proofs of these claims are deferred to Appendix B.1.3. Denote $\Delta_{1}=560229,\Delta_{2}=459951$, combining the claims gives us for all $f\in\ell_{2,\nu_{n}}$ that
$\operatorname{Var}_{\tau}(f)=T_{1}+T_{2}\leq\left(\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}\right)\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$
$=\frac{\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}}{\nu_{n}(K)}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}\nu_{n}(k)\mathcal{L}(k,k+1)$
$\leq\frac{\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}}{\nu_{n}(K)}\sum_{k=0}^{\infty}\left(f(k)-f(k+1)\right)^{2}\nu_{n}(k)\mathcal{L}(k,k+1)=\frac{\Delta_{1}n^{2-4\alpha}+\Delta_{2}n^{3-6\alpha}}{\nu_{n}(K)}\langle f,-\mathcal{L}f\rangle_{\nu_{n}}.$
Hence proved. ∎
6. Conclusion and Future Work
@@
In this paper, we establish bounds on the Chi-squared distance between the finite-time and the steady-state queue length distribution for the $M/M/n$ queue. We demonstrate that the mixing rate is tight (possibly up to a constant) in the many-server-heavy-traffic regimes. We also obtain bounds on the mean, moments, and tail of the queue length for a finite time and finite $n$. To prove these results, we build a Lyapunov-Poincaré framework and also propose two different ways to obtain the local Poincaré inequality which are of independent interest. In particular, our proposed local canonical path method and the truncation method have the potential to provide stronger bounds than the classical drift and minorization method (Down et al., 1995; Rosenthal, 1995).
@@
While we focus on analyzing a birth-and-death chain in this paper, note that our Lyapunov - Poincaré methodology is flexible enough to be applied to any reversible CTMC. Generalizing our results to other reversible Markov chains would pose the main challenge of constructing a suitable Lyapunov function resulting in suitable negative drift outside a suitable finite set. Nonetheless, the results in Section 3 are versatile, and one can hope to employ this methodology to obtain mixing results beyond birth-and-death chains. Beyond queueing theory, our finite set approaches can be further applied to other areas such as Learning, Generative AI and Sampling (Raginsky et al., 2017; Vempala and Wibisono, 2022; Durmus and Moulines, 2014; Chervonyi et al., 2025; Chen et al., 2024).
Acknowledgements.

### v4

@@ 5.2.3 Proof of Lemma 2
$\tau(q)$ $\geq\left(1+\frac{1}{4(n^{\alpha}-1)}\right)\frac{\left(1-\frac{1}{(n^{\alpha}-2)^{2}}\right)^{n}\sqrt{1-2n^{-\alpha}-n^{-1}}}{e^{\frac{1}{12\lfloor n-2n^{1-\alpha}\rfloor}}\left(2+n^{\alpha-1}\right)}\text{ for }q=n-1,$ (45)
$\tau(q)$ $\geq\frac{e^{\frac{n^{1-2\alpha}}{1-n^{-\alpha}}}}{1-n^{-\alpha}}\frac{\left(1-\frac{1}{(n^{\alpha}-2)^{2}}\right)^{n}\sqrt{1-2n^{-\alpha}-n^{-1}}}{e^{\frac{1}{12\lfloor n-2n^{1-\alpha}\rfloor}}\left(2+n^{\alpha-1}\right)}n^{-\alpha}\text{ for }q\in K-\{n-1\}.$ (46)
@@
Let $g_{1}:\mathbb{N}\rightarrow\mathbb{R}^{+}$ be a function to be defined later, we have $T_{1}\leq g_{1}(n)\times n^{2-4\alpha}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$. Here, we have $g_{1}(n)\leq 384\forall n\geq n_{0}$ and
$\lim_{n\rightarrow\infty}g_{1}(n)$ =2 $=2\text{ for }\alpha>1$/2, (47)
$\lim_{n\rightarrow\infty}g_{1}(n)$ $=2e^{3}\text{ for }\alpha=1/2.$ (48)
@@
Let $g_{2}:\mathbb{N}\rightarrow\mathbb{R}^{+}$ be a function to be defined later, we have $T_{2}\leq g_{2}(n)\times n^{3-6\alpha}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$. Here, we have $g_{2}(n)\leq 395.93\,\forall n\geq n_{0}$ and
$\lim_{n\rightarrow\infty}g_{2}(n)$ $=0\text{ for }\alpha>1/2,$ (49)
$\lim_{n\rightarrow\infty}g_{2}(n)$ $=e^{5}\text{ for }\alpha=1/2.$ (50)
@@
The full proofs of these claims can be found in Appendix 9.1.3. Combining the claims gives us
$\operatorname{Var}_{\tau}(f)=T_{1}+T_{2}$ $\leq\left(g_{1}(n)n^{2-4\alpha}+g_{2}(n)n^{3-6\alpha}\right)\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}Q_{K}(k,k+1)$
$=\frac{g_{1}(n)n^{2-4\alpha}+g_{2}(n)n^{3-6\alpha}}{\nu_{n}(K)}\sum_{k\in K^{\prime}}\left(f(k)-f(k+1)\right)^{2}\nu_{n}(k)\mathcal{L}(k,k+1)$
$\leq\frac{g_{1}(n)n^{2-4\alpha}+g_{2}(n)n^{3-6\alpha}}{\nu_{n}(K)}\sum_{k=0}^{\infty}\left(f(k)-f(k+1)\right)^{2}\nu_{n}(k)\mathcal{L}(k,k+1)$
$=\frac{g_{1}(n)n^{2-4\alpha}+g_{2}(n)n^{3-6\alpha}}{\nu_{n}(K)}\langle f,-\mathcal{L}f\rangle_{\nu_{n}}$
holds for all $f\in\ell_{2,\nu_{n}}$. Hence proved. $\square$
6 Conclusion and Future Work
@@
In this paper, we establish bounds on the Chi-squared distance between the finite-time and the steady-state queue length distribution for the $M/M/n$ queue. We demonstrate that the mixing rate is tight (possibly up to a constant) in the many-server-heavy-traffic regimes. We also obtain bounds on the mean, moments, and tail of the queue length for a finite time and finite $n$. To prove these results, we build a Lyapunov-Poincaré framework and also propose two different ways to obtain the local Poincaré inequality which are of independent interest. In particular, our proposed local canonical path method and the truncation method have the potential to provide stronger bounds than the classical drift and minorization method [23, 68].
@@
While we focus on analyzing a birth and death chain in this paper, note that our Lyapunov - Poincaré methodology is flexible enough to be applied to any reversible CTMC. Generalizing our results to other reversible Markov chains would pose the main challenge of constructing a suitable Lyapunov function resulting in suitable negative drift outside a suitable finite set. Nonetheless, the results in Section 3 are versatile, and one can hope to employ this methodology to obtain mixing results beyond birth and death chains. Beyond queueing theory, our finite set approaches can be further applied to other areas such as Learning, Generative AI and Sampling [66, 84, 26, 13].
7 Acknowledgement